OpenStudy (asdfghj8):

Solve the system of equations. *Will Medal* -5x+4y=-8 -9x+5y=1

OpenStudy (jango_in_dtown):

Method of cross multiplication is known to you?

OpenStudy (asdfghj8):

yes but isnt that just used for fractions? I know for this we have to eliminate one of the variables...

OpenStudy (jango_in_dtown):

well you can use one of the methods: 1st method: from the first equation find x in terms of y and then substitute this x value in the second equation

OpenStudy (jango_in_dtown):

then you will get a value of y..

OpenStudy (asdfghj8):

can you show me how to do that? every time i try i get it wrong

OpenStudy (jango_in_dtown):

Ok sure

OpenStudy (jango_in_dtown):

1st equation : -5x+4y=-8 so 5x=4y+8 implies x=(4/5)(y+2).

OpenStudy (jango_in_dtown):

wait.. lets do it in an easier way

OpenStudy (asdfghj8):

ok

OpenStudy (jango_in_dtown):

I am giving you a general rule for one step solution

OpenStudy (jango_in_dtown):

Suppose you have the following equations: ax+by+c=0 dx+ey+f=0

OpenStudy (jango_in_dtown):

then x/(bf-ec)=y/(cd-af)=1/(ae-db)

OpenStudy (asdfghj8):

ok

OpenStudy (jango_in_dtown):

first put the given equations in that form as I said

OpenStudy (jango_in_dtown):

then you have -5x+4y+8=0 -9x+5y-1=0

OpenStudy (asdfghj8):

i see that

OpenStudy (jango_in_dtown):

then use the general formula x/(-4-40)=y/(-72-5)=1/(-25+36)

OpenStudy (asdfghj8):

alrright

OpenStudy (jango_in_dtown):

so x=(-44/11)=-4 y=(-77/11)=-7

OpenStudy (asdfghj8):

so why dividing by 11?

OpenStudy (jango_in_dtown):

see what er did... check the coefficients a b c d e f

OpenStudy (asdfghj8):

ohhhhhhh ok

OpenStudy (jango_in_dtown):



OpenStudy (jango_in_dtown):

for y , omit the second column and do the same thing.